### v1

# SPAWNing Structural Priming Predictions from a Cognitively Motivated Parser

Abstract
Structural priming is a widely used psycholinguistic paradigm to study human sentence representations. In this work we propose a framework for using empirical priming patterns to build a theory characterizing the structural representations humans construct when processing sentences. This framework uses a new cognitively motivated parser, SPAWN, to generate quantitative priming predictions from theoretical syntax and evaluate these predictions with empirical human behavior. As a case study, we apply this framework to study reduced relative clause representations in English. We use SPAWN to generate priming predictions from two theoretical accounts which make different assumptions about the structure of relative clauses. We find that the predictions from only one of these theories (Participial-Phase) align with empirical priming patterns, thus highlighting which assumptions about relative clause better capture human sentence representations.
1 Introduction
[…]
Prior work found that targets like 1 were produced more often, and were processed more rapidly, when they were preceded by primes like 1, that have the same structure, than when they were preceded by primes like 1, which, while describing the same transfer event as 1, have a different structure.
. The lawyer sent the letter to the client.
. The lawyer sent the client the letter.
From this result, Branigan et al. (1995) inferred that participants’ mental representation of 1 is more similar to that of 1 than of 1 .
Branigan and Pickering (2017) propose that by carefully studying which sentences prime each other we can build a theory of human structural representations. Building such a theory requires us to generate hypotheses about what are the interesting prime-target pairs to compare. Insights from theoretical syntax, a field that has spent decades studying the structure of sentences, can help constrain this hypothesis space Gaston et al. (2017). In this work we introduce a new parser, the Serial Parser in ACT-R With Null elements (SPAWN), that can generate quantitative priming predictions from theories in syntax. By comparing SPAWN predictions from competing theoretical accounts, we can study which theoretical differences result in differing priming predictions, and therefore might be meaningful for psycholinguistics. Then, by comparing these predictions to empirical priming behavior in humans, we can study which theoretical assumptions are consistent with the representations that humans construct when processing sentences.
SPAWN is a cognitively motivated parser in which all of the parsing decisions are based on the computational principles proposed by a general purpose cognitive architecture, Adaptive Control of Thought-Rational (ACT-R; Anderson et al., 2004). Thus, SPAWN not only describes the computations underlying human parsing (Marr’s computational level), but also specifies the cognitive processes involved (Marr’s algorithmic level). This level of specification is necessary to explain why a sentence A is primed more by sentence B compared to C, and therefore necessary to generate quantitative behavioral priming predictions from syntactic theories. Existing algorithmic models of parsing Lewis and Vasishth (2005) can account for only a limited set of linguistic phenomena and theoretical assumptions. SPAWN can model a wider range of phenomena and assumptions because it uses a more flexible grammar formalism and specifies mechanisms to handle covert lexical items.
[…]
. The cat which was examined by the doctor was skittish.
We generate priming predictions from two competing syntactic theories: Whiz-Deletion (Chomsky, 1965), which assumes that the structure of 1 is identical to the structure of 1, and Participial-Phase (Harwood, 2018) which assumes that 1 and 1 have different structures. We describe these theories in more detail in § 4. We generate six sets of predictions from the two theories by modulating the strength of prior knowledge and the specific re-analysis mechanism used. Then, we compare the predictions from these two theories to empirical human data we collected using a novel web-based comprehension-to-production priming paradigm.
We found that the predictions from the Whiz-Deletion never aligned with the qualitative pattern of human priming behavior, whereas the predictions from Participial-Phase theory aligned with the empirical pattern when the predictions are generated from a SPAWN model with very weak to no prior knowledge; this observation held for the two variations of the reanalysis the mechanisms we tried. Taken together, these results suggest that the Participial-Phase account better characterizes human sentence representations, and more broadly highlight how SPAWN can be used to adjudicate between competing theoretical assumptions.
2 Background
2.1 ACT-R framework
ACT-R is a cognitive architecture with modules which are designed to explain general cognition through a small set of general computational principles and mechanisms that are relevant to a wide range of tasks and domains. One such mechanism which is particularly relevant in SPAWN is the retrieval of information from memory. The specific computational principles and algorithms that guide retrieval in ACT-R are outlined in § 3.2.1. Crucially, since ACT-R is intended to be a general purpose cognitive mechanism, most of the hyperparameters involved in this algorithm are already fixed based on data from a wide range of experimental paradigms and cognitive phenomena. This is useful because it restricts the degrees of freedom, thereby constraining the space of predictions that can be generated from any given theory.
2.2 Prior models of parsing
In most existing symbolic and neural network based parsers, the parsing decisions are not driven by specific cognitive principles such as the ones proposed by ACT-R. Therefore, generating predictions about observable human behavior (e.g., reading times) from these parsers requires making some additional linking hypotheses. Most prior hypotheses that link parsing decisions to human behavior have focused on notions of processing effort, such as the number of parse states explored Hale (2011), the maximum number of items on the stack at any given point Joshi (1990), or the maximum amount of time a node stays in memory De Santo (2021). These hypotheses cannot be used to generate priming predictions because they do not specify a mechanism by which a prime sentence might facilitate the processing of a target sentence.
One notable exception is the ACT-R based left-corner repair parser proposed by Lewis and Vasishth (2005), in which parsing decisions are made based on the activation of different chunks in the memory (such as words or grammar rules). The activation of chunks in this model can capture notions of both processing difficulty and priming. However, this model assumes a strong dissociation between the grammar and the lexicon and therefore cannot be adopted directly to generate predictions from lexicalized grammar formalisms such as Minimalist Grammar Stabler (1996), Combinatorial grammar Steedman (1988), Lexical-Functional Grammar Kaplan et al. (1981) or Head-Driven Phrase Structure Grammar Pollard and Sag (1994). The goal of this work is to develop a framework that can be used to generate priming predictions from any theory of syntax that can generate parse trees.
2.3 Prior models of priming
Reitter et al. (2011) proposed an ACT-R based model of priming in which lexical and grammatical knowledge are more strongly interconnected than in the parser proposed by Lewis and Vasishth (2005). However, this model can only generate sentences given a semantic description, and therefore can only be used to model sentence production and not sentence processing. The models of priming that can model processing either do not explicitly model syntactic structure (Chang et al., 2006; Malhotra, 2009; Prasad et al., 2019; Sinclair et al., 2022) or do not explicitly implement the priming mechanism (Snider, 2008). Hence, the goal of this work is to bridge this gap and develop a model that explicitly models both syntactic structure and the priming mechanism, which can be used to generate behavioral processing predictions.
3 Model description
SPAWN uses the three components of ACT-R that are relevant for parsing: declarative memory (§ 3.1), which contains information about lexical and syntactic categories (cf. Reitter et al., 2011); procedural memory (§ 3.2), which contains the algorithm for retrieving syntactic categories from memory and combining them together; and buffers, which store the words the parser has encountered so far, the syntactic categories retrieved for those words and the current parse state.
3.1 Declarative memory (the grammar)
Declarative memory in SPAWN consists of two types of chunks (sets of attribute-value pairs): syntax chunks and lexical chunks.
Lexical chunks:
Each lexical chunk stores a lexical item in the vocabulary along with the set of syntactic categories that the word could be associated with. For example, the lexical chunk for the word “by” encodes that it needs to be associated with the syntactic category “preposition”, whereas the lexical chunk for “examined” encodes that it can be either be associated with the “transitive verb” category or the “past participle” category.
Syntax chunks:
Each syntax chunk stores the constraints on the contexts in which a category can occur. For example, the past participle category encodes that it needs to have a prepositional phrase category on its right; transitive verb encodes that it needs to have a determiner phrase category on its left and right.111The CCG notation to encode the “transitive verb” category is (TPDP)/DP; the forward slask indicates the words needs to combine with a DP on the right and the forward slash that it needs to combine with DP on the left. TP is the category that results from this combination. Since Combinatorial Categorial Grammar (CCG; Steedman, 1988) is designed to express such constraints, we use this grammar formalism to represent our syntax chunks.
3.2 Procedural memory (the parser)
SPAWN parses sentences incrementally, one token at a time. For each token, it retrieves a syntactic category (§ 3.2.1), and tries to integrate it with the current parse state (§ 3.2.2). If integration is not successful, then it retrieves another category and repeats the process. If none of the possible tags can successfully integrate with the current parse state, a reanalysis is triggered; depending on the model setting, the parser either goes back to the start of the sentence, or selects a previous token to reanalyze (§ 3.2.3). Once the integration is successful, the parser decides whether the next token is a null token (a token not associated with any overt word in the sentence) and if yes, adds it to the sequence of tokens (§ 3.2.4). Finally, the parser proceeds to the next token. This retrieval-integration-reanalysis-null prediction continues until all of the tokens in the sentence are assigned syntactic categories.
3.2.1 Retrieval
When processing a word , the parser retrieves the category with the highest activation from the set of all possible categories that can be associated with. The activation for any category is given by Equation 1, where is the base-level activation, is the activation spreads to , is the inhibition from the buffer to the when processing sentence , and is noise that is sampled from a specified normal distribution.
Base-level activation:
This activation for a category is high if the category has been retrieved recently and/or frequently. It is given by Equation 2, where is the total number of times the model has encountered , is the time taken to process all the words since the model’s -th encounter of , and is a decay parameter.
The time to process a word is given by Equation 3, where is the number of chunks retrieved when processing , the activation of the -th chunk the model retrieved when processing (as computed using Equation 1), a latency factor and a latency exponent; these hyperparameters are set based on prior ACT-R models (see Appendix 5.3).
Lexical activation:
The context independent activation a word spreads to a category is given in Equation 4, where , another hyperparameter set based on prior ACT-R models (see Appendix 5.3), is the maximum activation that any word can spread.
Inhibition:
The inhibition for takes into account how often was retrieved for but was later discarded during reanalysis when processing a sentence . It increases if was discarded often and/or recently, and is given by Equation 5 where indicates the total number of times was discarded when processing in the current sentence, indicates the time since the -th time was discarded in sentence , and is the decay factor.
3.2.2 Integration
Integrating a retrieved syntactic category involves combining with the current parse state ; this combination is determined by the CCG composition process (Steedman, 1996). The parser starts by sequentially trying to apply each of the six composition rules (see Appendix B.2), stopping once a successful combination is found. If no successful combination is found, then the parser tries to apply each of the type-raising rules222A type-raising rule allows the incremental composition of categories. For example consider the category for transitive verb (TPDP)/DP. The parentheses dictate that the DP on the right of the verb needs to combine with the verb before the DP on the left. This is a problem because words are processed from left-to-right in English. A type-raising rule raises the DP on the left of the verb to TP/(TPDP), which lets it combine with the transitive verb category to result in the state TP/DP. in the grammar to the retrieved tag or parse state and then sequentially apply all six composition rules. If no successful combination is possible after type raising, then the retrieved category cannot be integrated into the current parse state; the parser then needs to either retrieve another category for the word being processed, or, if no unexplored categories remain, trigger a reanalysis.
3.2.3 Reanalysis
SPAWN uses an inhibition-based backtracking mechanism for reanalysis. When a reanalysis gets triggered at , the parser selects an index to backtrack to, where . We use two methods to select the reanalysis index. In one, we always select the first word of the sentence. In the other, we sample the index weighted by the number of categories that words at each index can be associated with (i.e., the amount of context-independent uncertainty at each index). Once it selects , it discards all of the categories retrieved for , and resets the parse state to what it was at . The parser keeps track of the categories that were discarded for all the words when processing a given sentence , and uses this to compute the inhibition for each category using Equation 5.333While storing all of the discarded categories is a convenient way to compute inhibition, it is not cognitively plausible. Future work can examine other ways of computing inhibition that relies on summaries of discarded categories, instead of storing all of the instances.
While the inhibition term decreases the probability of repeatedly retrieving the same incorrect category for a word every time it is reanalyzed, it does not eliminate it: inhibition might be insufficient to suppress the relative activation of an incorrect category , it has very high base-level or lexical activation. To ensure that the parsing process eventually results in the right parse, we implemented a “give-up” mechanism, where after iterations (set to ), the model uses only inhibition and noise to compute activation of .
3.2.4 Null or covert elements
SPAWN assumes that only certain parse states can be followed by null elements; this assumption ensures that the parser does not try to insert null elements after every token, which would result in an exponential increase in the search space because every null token could be followed by other null tokens. When the current parse state can be followed by a null element, there is uncertainty about the next parsing decision. Consider the following sentence, which is not related to relative clauses or part of our experimental setup, but illustrates the uncertainty in a theory-independent way.
. The cat examined the doctor and the doctor examined the cat.
This sentence is grammatical with or without the bolded red word. Therefore, after parsing the prefix “The cat examined the doctor and the doctor”, the parser has to predict whether the upcoming verb in the sentence is covert or overt. If it expects the next verb to be a covert “examined”, it should retrieve a null transitive verb category and integrate it with the current parse state before processing the remainder of the sentence. On the other hand, if it expects an overt “examined”, the parser should not retrieve any null elements — retrieving a null transitive verb would make the sentence ungrammatical when it encounters the overt transitive verb.
We model this decision in SPAWN in the same way that we model other uncertainty: pick the option with the highest activation, where corresponds to the current word and , where correspond to the different types of null elements that can come after the current parse state. The activation for is given by Equation 6:
is computed the same as lexical activation for syntactic categories (Equation 4), and the same as inhibition (Equation 5). As in Equation 1, is noise sampled from a normal distribution. We do not include base-level activation for the null categories in this computation, because the base-level activation for the not-null category would be extremely high — most sentences in the corpus do not have null elements —and would therefore dominate the overall activation, resulting in the null categories never being retrieved.
3.3 Updating activations (“learning”)
Learning in SPAWN occurs by updating the counts of syntactic categories, which in turn are used to compute base-level and lexical activations (Equations 2, 4). These counts are updated at the end of processing each sentence based on the final set of categories and null-elements that were retrieved.
3.3.1 How does priming emerge?
SPAWN does not explicitly model priming. Rather, priming emerges as a consequence of parsing and learning. There are two reasons priming emerges: an increase in the activation of relevant categories and an increase in the probability of reanalysis.
Increased activation:
[…]
. The lawyer sent the client the letter.
Let us assume that in 3.3.1, the verb “sent” is associated with a PO category, whereas in 3.3.1 it is associated with a DO category. Therefore the base-level activation for PO category, as well as the lexical activation from “sent” to PO will be higher in a model that has just parsed 3.3.1 than in a model that has just parsed 3.3.1. Now consider that a model which has processed one of these sentences is presented with a partial target sentence “The boy sent the ___”. The model is more likely to retrieve the PO category for “sent”, and thus more likely to produce a sentence like 3.3.1 if it has just processed 3.3.1, compared to if it has just processed 3.3.1.
. The boy sent the gift to his friend.
Increased probability of re-analysis:
Consider a model that just processed a sentence like 3.3.1, where the verb “sent” is assigned SHIFT category.
. The lawyer sent to his client the letter.
After processing this sentence, the model is more likely to expect a “to” after the verb “sent” than “the”. Therefore, when it encounters the partial target “The boy sent the” a reanalysis gets triggered. The probability of the model assigning PO category to “sent” is higher when it initially retrieves a SHIFT category and a reanalysis gets triggered, than when it initially retrieves a DO category. Therefore, SPAWN would predict that the probability of the target being completed with 3.3.1 is higher when it is preceded by 3.3.1 compared to 3.3.1.
4 A case study: competing theories of reduced relative clauses
We use SPAWN to generate and test priming predictions for two competing syntactic theories of relative clauses that differ in their assumptions about how the structure of sentences like 4 is related to the structure of sentences like 4 and 4.
. The cat examined by the doctor was skittish. (Reduced passive RC; RRC)
. The cat who was examined by the doctor was skittish. (Full passive RC; FRC)
. The cat being examined by the doctor was skittish. (Reduced progressive RC; ProgRRC)
The Whiz-Deletion account of RCs (Chomsky, 1965) argues that the sub-tree corresponding to any RC, whether it is reduced or not, is headed by the same node: a complementizer phrase (CP). In full RCs, the lexical content in this phrase (the wh-word and auxiliary “was”) is overt, whereas in reduced RCs this lexical content is covert. The Participial-Phase account (Harwood, 2018) argues that full RCs are headed by CPs, and therefore include the wh-word and auxiliary “was”. Reduced passive and progressive RCs on the other hand are headed by Voice Phrase (VoiceP) and Progressive Phrase (ProgP) respectively. These nodes are also present in full RCs, but are nested deeper inside the CP. Consequently, the Participial-Phrase account, unlike the Whiz-Deletion account, does not assume the presence of a covert wh-word and auxiliary in reduced passive and progressive RCs. See § A.1 for trees that illustrate these differences.
Implementing the two theories
The differences between the two theories are implemented in the declarative memory (or grammar), summarized in § A.2; the procedural memory (parsing mechanism) is identical across both theories.
The differences in assumptions about the head of the RC are represented by different categories for nouns that are modified by RCs. In the Whiz-Deletion grammar, all nouns modified by RCs have the category NP/CP (i.e., a noun looking to combine with a CP on its right). In the Participial-Phrase grammar, nouns modified by FRCs have the NP/CP category, nouns modified by RRCs have the NP/VoiceP category, and nouns modified by ProgRRCs the NP/(VoiceP/ProgP) category.
The differences in assumptions about the presence of covert lexical items is represented by different null categories. For RCs which modify the subject of the clause, the Whiz-Deletion grammar has categories associated with the null wh-word, CP/(TPDP), and with the null auxiliary, (TPDP)/VoiceP, which when combined together result in the parse state CP/VoiceP, i.e., a state looking for a passive continuation. Both of these null categories are absent from the Participial-Phase grammar.444Both accounts have the same categories for the null wh-word in RCs which modify objects of clauses like “The cat the doctor examined was skittish”.We also experimented with a version of the Whiz-Deletion grammar that combines the two null categories into one, but found that this results in nearly identical predictions.
5 Methods
5.1 Priming Experimental paradigm
[…]
. The monkey chased the hatter and stole a hat.
. The dentist chased her son and panted.
. The thief chased ___
5.2 Estimating priming effects from humans
Participants
We recruited 769 US-based participants from Prolific, of whom 765 were self-reported native speakers of English. We compensated them with 8.35 USD.
Design
We developed a web-based version of the comprehension-to-production priming paradigm used by Pickering and Branigan (1998). In the original paradigm, participants were given incomplete sentences in a booklet and asked to complete them. Since participants can be less attentive on web-based platforms than in the lab, we modified the paradigm to ensure that participants had to fully read the prime sentences. On the prime trials, participants were presented with a sentence, and asked to re-type that sentence from memory on the next screen. They could not progress until they typed in the sentence perfectly, and could not copy-paste the sentence, but could go back to re-read the sentence as often as they liked. On the target trials, participants were presented with the partial prompt on the screen, and asked to re-type the prompt and complete it on the next screen. They could not progress until they typed in the prompt perfectly and entered at least one more word. We did not automatically verify participants’ productions, but in practice almost all participants generated grammatical completions with real words.
Materials
[…]
. The dog who was chased by the boy ran away.
. The dog being chased by the boy ran away.
. The dog chased the boy and ran away.
From these materials we created counterbalanced lists: in each list, three items occurred as primes; the fourth was cut at the verb to generate the target.
Statistical model
We used regular expressions to classify all target completions into two categories: RRC vs. non-RRC. To estimate , we fit a Bayesian mixed-effects logistic regression model with the following three predictors (specified using Helmert contrasts) as fixed effects: All RCs vs. AMV, ProgRRC and FRC vs. RRC, and ProgRRC vs. FRC. We used a weakly informative prior and a maximal random effects structure with by-item and by-participant random intercepts and slopes (see Appendix D).
5.3 Generating priming predictions
We generated predictions from 12 types of models which varied along 3 dimensions: the grammar (Whiz-Deletion vs. Participial-Phase), amount of training data (0, 100, 500 sentences), and type of reanalysis (back to start vs. sample based on uncertainty). For each type, we created 1280 model instances by sampling different hyperparameters.
Model hyperparameters
There are two types of parameters in SPAWN. First, parameters that are fixed across all instances: decay ( in Equations 2, 5) set to , latency exponent ( in Equation 3) set to and maximum activation ( in Equation 4) set to . Second, parameters that differ for each model instance: latency factor ( in Equation 3) sampled from , and the noise parameter (the standard deviation of the normal distribution used to sample in Equation 1) sampled from . All of the parameters or the distributions they were sampled from, except for the noise parameter, were taken from Vasishth and Engelmann (2021). We sampled the noise parameter from instead of used in prior work because initial explorations revealed that when the noise parameter was sampled from this constrained range, models almost never retrieved syntactic categories with low base-level activation.
Training data
To approximate the base-level and lexical activations that participants start with prior to the experiment, we trained SPAWN instances on either 100 or 500 sentences. These numbers are consistent with prior work which assume that participants start experiments with very weak priors (Delaney-Busch et al., 2019; Fine et al., 2010). We used templates to generate a dataset of 10000 sentences in which the relative frequency of different types of RC sentences mirrored corpus statistics from Roland et al. (2007); for example, only 1% of the training sentences contained an RRC (see Appendix § C) For each model instance, we sampled the training sentences from this dataset. Given the low probability of RRCs, many model instances never encountered these in their training data, and started with a base-level activation of for RRCs.
Priming experiment
We used the same stimuli lists from the human experiment to measure priming effects in the model instances, and the same statistical model to compute . On each prime trial, the model parsed the sentence and updated the base-level and lexical activations based on the final set of retrieved categories. On each target trial, the model parsed the partial prompt and we recorded the resulting parse state; the model was constrained to end with only one of two partial states: DP/PP (RRC parse) or TP/DP (active parse).555Other partial parse states would require completions with complex syntactic structures that people do not produce.
6 Results
Participant/model exclusion
Most participants (77%) never generated a single RRC target completion. Similarly most models (median of 66% across the 12 model types) never generated a single RRC parse state. Since the goal of this work is to find differences in the proportion of RRC completions across the primes, we only included in our statistical analyses the participants or models that generated at least one RRC completion or parse.
Participial-Phase vs. Humans
The Participial-Phase model instances trained on no data with both re-analysis mechanisms captured the empirical priming pattern qualitatively; they also perfectly captured the magnitude of the priming effect in the RRC condition, and approximately captured the magnitude in the other conditions. With an increase in training data, the priming predictions diverged from the empirical data in two ways: First, the overall magnitude of priming decreased. Second, the relative differences between the priming conditions changed; this was especially salient in model instances which sampled the word to reanalyze based on uncertainty (as opposed to going back to the start) where the other RC conditions patterned more closely with the non-RC condition.
Whiz-Deletion vs. Humans
None of the Whiz-Deletion models captured the qualitative priming pattern observed in humans: the predicted effect for ProgRRC was higher than for FRC; empirically the priming effect in the two conditions was equal. The magnitude of priming in the RRC condition was also smaller than what was observed empirically.
7 Discussion
In this work we introduced a cognitively motivated parser, SPAWN, which can be used to generate quantitative behavioral predictions from any theory of syntax capable of generating incremental parse trees. SPAWN makes it possible to evaluate what theoretical differences (if any) result in differing sentence processing predictions. As a case study, we used SPAWN to generate predictions from two competing theories of reduced relative clauses (Whiz-Deletion and Participial-Phase) while modulating the reanalysis mechanism (going back to the start vs. sampling based on uncertainty) and the number of training examples (0, 100, 500). We compared the predictions from these different versions of the SPAWN model to human behavior from a large-scaled (N=769) web-based comprehension-to-production priming experiment.
We found that the predictions from the Whiz-Deletion SPAWN models did not capture the qualitative human priming behavior for any of the different reanalysis mechanisms or training sizes. In contrast, the Participial-Phase SPAWN models captured the human priming pattern, but only when the model was untrained (i.e., had no prior knowledge); As the prior knowledge increased, the relative differences between the priming conditions in humans and Participial-Phase models deviated, reinforcing assumptions from prior work that participants enter the experiment with weak prior knowledge. Taken together, these results suggest that the Participial-Phase account of reduced relative clauses captures the the structural representations people construct better than the Whiz-Deletion account.
Conclusion
We proposed a framework to generate priming predictions from syntactic theory. Using relative clauses as a case study, we demonstrated how this framework can be used to adjudicate between competing syntactic theories and parsing mechanisms.
8 Limitations
Like all computational models, SPAWN makes many simplifying assumptions, such as assumptions about serial parsing or the assumptions underlying the specific mechanisms we implemented for reanalysis, “giving up” and prediction of null elements. While we explored the consequences of some of these assumptions when we modulated the reanalysis mechanism, this exploration was not exhaustive. It is important for future work to more systematically study the impact that the different simplifying assumptions have on behavioral predictions, which in turn can suggest concrete ways of improving the model. Since SPAWN is already capable of generating more fine grained predictions about the time take to read or reprocess speciic words in sentences — the time taken for any of the parsing steps (retrieval, integration and reanalysis) is measured in milliseconds by default in ACT-R — a promising direction for such a systematic study would be to evaluate these fine-grained predictions from SPAWN with empirical eye-tracking data. In a similar vein, applying this framework to other empirical phenomena can also shed light on how to improve the model.
Additionally, the current version of SPAWN is implemented using the CCG grammar formalism. Therefore, to use SPAWN to generate predictions from some theory, it is necessary to implement the structural assumptions in that theory using CCG notation. While CCG is a very flexible formalism that can be used to express the assumptions of most syntactic theories — for example, in this work we converted Minimalist Grammar trees to CCG trees — future versions of SPAWN can add support for other grammar formalisms to minimize the need for translating between grammar formalisms.
Appendix A Details about the two theories of reduced relative clauses
A.1 Syntax trees
A.2 Differences in syntactic categories
| Category | Example sentence | Whiz-Deletion | Participial-Phase |
| Noun | “The cat which was examined by… ” | NP/CP | NP/CP |
| (“cat”) | “The cat examined by … ” | NP/CP | NP/VoiceP |
| “The cat being examined by … ” | NP/CP | NP/(VoiceP/ProgP) | |
| Null wh subject | “The cat NULLwh NULLaux examined by…” | CP/(TPDP) | MISSING |
| Null finite auxiliary | (TPDP)/VoiceP | MISSING | |
| Null progressive auxiliary | ‘The cat NULLwh NULLaux being examined by…” | (TPDP)/(VoiceP/ProgP) | MISSING |
Appendix B SPAWN parsing details
B.1 Schematic of all the components
B.2 CCG combination and type-raising rules
| Rule name | Parser state form | Tag form | Composed form |
| Forward composition | DP/NP | NP | DP |
| Backward composition | DP | TPDP | TP |
| Forward harmonic composition | DP/VoiceP | VoiceP/PP | DP/PP |
| Backward harmonic composition | TPDP | eosTP | eosDP |
| Forward crossed composition | CP/TP | TPDP | CPDP |
| Backward crossed composition | TP/VoiceP | eosTP | eos/VoiceP |
We have just one type-raising rule: DP can get type-raised to TP/(TPDP). This lets the subject DP in a sentence combine with a transitive verb — (TPDP)/DP — before the transitive verb combines with the object DP.
Appendix C Details about the training dataset
| Structure | Prob | Example |
| Subject RC | 0.016 | The defendant who examined the lawyer … |
| Full object RC | 0.002 | The defendant who the lawyer examined … |
| Reduced object RC | 0.005 | The defendant the lawyer examined … |
| Full passive RC | 0.002 | The defendant who was examined by the lawyer … |
| Reduced passive RC | 0.011 | The defendant examined by the lawyer … |
| Full progressive RC | 0.0002 | The defendant who was being examined by the lawyer … |
| Reduced progressive RC | 0.005 | The defendant being examined by the lawyer … |
| Transitive NP object | 0.321 | The examined the lawyer. |
| Transitive PP object | 0.080 | The defendant went to the store. |
| Intransitive | 0.240 | The defendant sang (joyfully). |
| Copular | 0.240 | The defendant was happy. |
| Coordination | 0.080 | The defendant examined the lawyer and went to the store. |
| The defendant was happy and sang joyfully. | ||
| The defendant went to the store and sang and was happy and examined the lawyer. |
Appendix D Details about statistical models
D.1 Model specification
To generate quantitative predictions about the predicted proportion of passive responses while taking into consideration the model-instance wise and item wise variation, we fit Bayesian mixed effects logistic regression models. We used a Helmert contrast coding scheme with the following predictors, which let us evaluate if the mean log odds ratio of the ProgRRC and FRC conditions are equal to each other and to the mean log odds ratio of the RRC condition.
C1: Compare the mean log odds ratio of the AMV condition with the mean log odds ratio of all the RC conditions combined.
C2: Compare the mean log odds ratio of the RRC condition with the mean log odds ratio of all ProgRRC and FRC conditions combined.
C3: Compare the mean log odds ratio of the ProgRRC condition to the mean log odds ratio of the FRC condition.
We fit the maximal model by including all by-model-instance and by-item random intercepts and slopes.
D.2 Priors
we fit the model using the following weakly informative prior.
| Intercept | |||
| Fixed effects | |||
| SD for random effects |
This prior assumes that the log odds ratio between priming conditions is most likely to be 0 (i.e. no priming effect) and unlikely to be greater than 4 or less than -4. This assumption is based on a meta-analysis of priming in production studies (mahowald2017) where the log odds ratio between the prime conditions was not greater than 4 in any of the constructions they considered.
D.3 Statistical inferences
From the posteriors of the Bayesian models, we computed 95% credible intervals for P(RRC prime, target) for each prime condition and model type. We also computed these 95% for human data. We use the overlap in these credible intervals to draw our statistical inferences. If the credible intervals in two priming conditions are non overlapping, we infer that the priming effect in these conditions are different. Similarly, if the credible intervals for predicted priming effect from a model does not overlap with the empirical priming effects, we infer that the model cannot account for human behavior. Such an inference is valid because credible intervals, unlike the frequentist confidence intervals, reflect our confidence about the distribution of the actual effects (so 95% credible interval means that we are 95% sure that the true effect falls within this interval).